OpenStudy (anonymous):

LaTex....how do I show the evaluation of an integral? \[|x^2 \left_{-2}^{14}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

not too concerned about the numbers...but how would I write that in Latex?

OpenStudy (zarkon):

\[\left.\int\limits_{a}^{b}f(x)dx=F(x)\right|_{a}^{b}=F(b)-F(a)\]

OpenStudy (turingtest):

\int_a^b f(x)dx=\left.F(x)\right]_a^b \[\int_a^b f(x)dx=\left.F(x)\right]_a^b\] is what I use

OpenStudy (zarkon):

aslo..post questions like this here http://openstudy.com/study#/groups/LaTeX%20Practicing!%20%3A)

OpenStudy (anonymous):

\[x^2\bigg|_2^3\]

OpenStudy (anonymous):

\bigg|

OpenStudy (zarkon):

that is a nice way to do it

OpenStudy (anonymous):

yes...thanks again everyone!

OpenStudy (anonymous):

@MathSofiya anything about latex you can find easily by googling

OpenStudy (zarkon):

\Bigg| will make it just a little longer...but not much